OpenStudy (girlgoyleh.):

Please help, will medal and fan!!! In a particular year, the population of Dallas, TX, was approximately 1 x 10^6 people. If the population of Los Angeles, CA, was approximately three times that of Dallas, there were about ? x?^ 6 people living in Los Angeles that year. The population of Pittsburgh, PA, was approximately 3 x 10^5 people during the same year, which was about ? times ?that of Dallas.

OpenStudy (anonymous):

1,00,00 i would think

OpenStudy (aum):

"population of Los Angeles, CA, was approximately three times that of Dallas" Population of Los Angeles = 3 x 1 x 10^6 = 3 x 10^6

OpenStudy (anonymous):

wait....

OpenStudy (anonymous):

yeah aum is right

OpenStudy (anonymous):

3 x 1 = 3 then it would be 3 x 10^6

OpenStudy (girlgoyleh.):

the last question mark is "more than or less than?"

OpenStudy (aum):

\[ \frac{\text{Population of Pittsburgh }}{\text{Population of Dallas}} = \frac{3 \times 10^5}{1 \times 10^6} = \frac{3}{10} = 0.3 \]

OpenStudy (girlgoyleh.):

so it is less than

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

:)